# DISCO: Distribution-Aware Calibration for Object Detection with Noisy Bounding Boxes

### v2

@@ Abstract
Large-scale well-annotated datasets are of great importance for training an effective object detector. However, obtaining accurate bounding box annotations is laborious and demanding. Unfortunately, the resultant noisy bounding boxes could cause corrupt supervision signals and thus diminish detection performance. Motivated by the observation that the real ground-truth is usually situated in the aggregation region of the proposals assigned to a noisy ground-truth, we propose DIStribution-aware CalibratiOn (DISCO) to model the spatial distribution of proposals for calibrating supervision signals. In DISCO, spatial distribution modeling is performed to statistically extract the potential locations of objects. Based on the modeled distribution, three distribution-aware techniques, i.e., distribution-aware proposal augmentation (DA-Aug), distribution-aware box refinement (DA-Ref), and distribution-aware confidence estimation (DA-Est), are developed to improve classification, localization, and interpretability, respectively. Extensive experiments on large-scale noisy image datasets (i.e., Pascal VOC and MS-COCO) demonstrate that DISCO can achieve state-of-the-art performance, especially at high noise levels.
@@ 1 Introduction
Object detection has made substantial progress in recent years Ren et al. (2015); Lin et al. (2017); Carion et al. (2020); Sun et al. (2021); Zou et al. (2023), which is largely attributed to the utilization of large-scale well-annotated datasets Everingham et al. (2010); Lin et al. (2014). However, obtaining accurate bounding box annotations is labor-intensive and demanding, especially for some real-world scenarios such as medical diagnosis Luo et al. (2021); Chai et al. (2023) and autonomous driving Michaelis et al. (2019); Mao et al. (2022). As shown in Figure 1(a), inherent ambiguities of bounding boxes are often caused by object occlusion or unclear boundaries He et al. (2019). Moreover, insufficient domain expertise and the strenuous workload can also lead to low-quality labeling of bounding boxes Liu et al. (2022). In practical deployments and applications, vanilla object detectors will inevitably suffer from such noisy bounding box annotations. Therefore, it is of scientific interest to explore how to tackle noisy bounding boxes in object detection.
Figure 1: (a) Examples of noisy bounding box annotations. For simplicity, the noisy annotation for one object is presented per image. (b) Trait comparison of existing solutions and our DISCO. Their learning behaviors for one single border of bounding boxes are presented above as an illustration. Note that our DISCO can achieve improvement in all three aspects.
Due to the degenerated supervision introduced by noisy annotations, object detection with noisy bounding boxes remains a challenging problem. Obviously, such corrupt supervision signals could weaken the localization precision of object detectors. Besides, although the classification accuracy is less affected Liu et al. (2022), noisy bounding boxes do introduce biased category features during training, which reduces the generalization capability of classification. Notably, there is also a significant concern about the lack of interpretability for box predictions, especially considering the influence of noisy bounding box annotations.
Encountering the above challenges, existing solutions still exhibit drawbacks in this special setting (see Figure 1(b)). Several previous works are dedicated to correcting noisy bounding boxes Liu et al. (2022); Li et al. (2020a); Xu et al. (2021), aiming to mitigate the effect of noisy box annotations. However, their performance gains in classification and localization are constrained by heuristic box correction approaches, and the detector cannot identify which bounding boxes are inaccurately predicted. An alternative series of methods focus on equipping the model with the capability to estimate the confidence of predicted bounding boxes He et al. (2019); Li et al. (2020b); Choi et al. (2019), through which the detector can be more robust against noisy box annotations. Despite these efforts, it is unfortunate that the detector is still plagued by flawed supervision during training.
@@
Figure 2: Examples of proposal aggregation in object detection with noisy bounding boxes. It can be observed that the real ground-truth is usually situated in the aggregation region of the proposals assigned to a noisy ground-truth.
@@
Essentially, the corrupt supervision signals should be blamed for the above issues. In this work, we expect to answer the following question: How to properly calibrate the corrupt supervision signals? As shown in Figure 2, we observed that the real ground-truth is usually situated in the aggregation region of the proposals assigned to a noisy ground-truth, showing that the spatial distribution of proposals can act as a statistical prior for the potential locations of objects. Thus, we propose DIStribution-aware CalibratiOn (DISCO), which aims to model the spatial distribution of proposals for calibrating supervision signals (see Figure 1(b)). For each group of the assigned proposals, we perform spatial distribution modeling with a four-dimensional Gaussian distribution, statistically extracting potential locations of objects. Based on the modeled distribution, we develop three distribution-aware techniques to improve classification, localization, and interpretability, respectively: 1) Distribution-aware proposal augmentation (DA-Aug): Additional proposals are generated from the distribution to enrich category features in the representative locations, and then the proposal with the highest classification score is collected to boost classification performance; 2) Distribution-aware box refinement (DA-Ref): With a non-linear weighting strategy, noisy ground-truth is fused with the distribution into a refined ground-truth to achieve superior bounding box regression; 3) Distribution-aware confidence estimation (DA-Est): An extra estimator is integrated into the detection head, with the distribution variance elegantly acting as its supervision, to estimate the confidence of predicted bounding boxes. Without introducing complicated learnable modules, DISCO can attain state-of-the-art performance on large-scale noisy image datasets (i.e., Pascal VOC and MS-COCO), especially at high noise levels. Our main contributions are summarized as follows:
  • Motivated by the observation about proposal aggregation, we propose an approach called DISCO to calibrate supervision signals with spatial distribution modeling.
  • To improve classification, localization, and interpretability, we introduce three techniques (i.e., DA-Aug, DA-Ref, and DA-Est) to collaborate with the modeled distribution in a distribution-aware manner.
  • Comprehensive experiments show that our DISCO can attain state-of-the-art performance in object detection with noisy bounding boxes and meanwhile achieve satisfactory interpretability for its predictions.
Figure 3: Training pipeline with our DISCO. DISCO is performed twice in a training iteration, the first time for proposal re-assignment to obtain better-matching proposals and the second time for producing calibrated supervision signals. In DISCO, spatial distribution modeling (Section 3.2) is performed firstly, followed by three distribution-aware techniques, i.e., DA-Aug (Section 3.3), DA-Ref (Section 3.4), and DA-Est (Section 3.5), to collaborate with the modeled distribution. Note that we additionally integrate an estimator into the vanilla detection head to construct the distribution-aware head (DA head) for the implementation of DISCO.
@@ 2.1 Object Detection
The goal of object detection is to recognize what objects are present and where they are situated. Faster-RCNN Ren et al. (2015) is a classic detection framework with a two-stage strategy, and is widely adopted and improved in subsequent works Cai and Vasconcelos (2018); Pang et al. (2019); Sun et al. (2021). Moreover, RetinaNet Lin et al. (2017), YOLO Redmon et al. (2016), and CenterNet Zhou et al. (2019) delve into strengthening the performance of one-stage detectors. Recently, transformer-based detectors Carion et al. (2020); Zhu et al. (2020); Zhang et al. (2022) also attract the attention of the community, which conducts object detection in an end-to-end fashion. Training with accurate bounding box annotations, object detectors can achieve satisfactory and even remarkable performance. However, object detection with noisy bounding boxes remains an under-explored subproblem.
2.2 Object Detection with Noisy Annotations
@@
Specifically, noisy annotations of an object detection dataset could compose of noisy category labels and noisy bounding boxes. Previous works Chadwick and Newman (2019); Li et al. (2020a); Xu et al. (2021) have made attempted to jointly tackle these two types of noisy annotations. Unlike this setting, we focus on training an object detector with noisy bounding boxes, since box noise is more common and challenging in realistic scenarios Liu et al. (2022). The state-of-the-art for this tough task is called OA-MIL Liu et al. (2022), which adopts a multi-instance learning (MIL) framework at the object level to correct bounding boxes. Besides, some approaches aiming to boost the robustness of detectors, such as KL Loss He et al. (2019), can also contribute to performance improvement in this task.
2.3 Weakly Supervised Object Detection
@@
Weakly supervised object detection (WSOD) is also a relevant task, where only image-level labels can be accessed to train an object detector. The mainstream solution is to treat WSOD as a MIL problem Bilen and Vedaldi (2016); Wan et al. (2018); Chen et al. (2020); Tang et al. (2017), where each training image is constructed as a bag of instances. To handle the non-convex optimization of MIL, spatial regularization Diba et al. (2017); Wan et al. (2018), optimization strategy Tang et al. (2017); Wan et al. (2019), and context information Kantorov et al. (2016); Wei et al. (2018) are introduced to attain better convergence. Moreover, it is worth noting that SD-LocNet Zhang et al. (2019a) contributes a self-directed optimization strategy to handle object instances with noisy initial locations. Unfortunately, WSOD always results in relatively inaccurate box predictions due to the lack of fine-grained supervision. Effective methods for object detection with noisy bounding boxes can contribute to further refining these box predictions.
@@ 3 Methodology
3.1 Overview
In this work, we propose DISCO to calibrate the corrupt supervision signals caused by noisy bounding boxes in object detection. Essentially, DISCO is a training-time calibration approach designed for two-stage detectors. In a training iteration, DISCO is performed twice using the distribution-aware head (DA head) with the assigned proposals as input (see Figure 3). These two times of DISCO follow the same process and have only subtle differences for different purposes. The first time aims to yield a refined ground-truth for proposal re-assignment, by which better-matching proposals can be obtained. The second time aims to produce spatial distributions of proposals acting as superior supervision. In the following, we start by describing spatial distribution modeling in Section 3.2. Then, we will introduce three distribution-ware techniques (i.e., DA-Aug, DA-Ref, and DA-Est) in Section 3.3, 3.4 and 3.5, respectively, in which we will also detail the differences between these two times of DISCO.
Figure 4: Illustration of spatial distribution modeling. For clarity, we present the process in the view of the whole bounding box and one single border. Note that the length of the vertical line indicates its weight. Here the refined ground-truth is essentially a spatial distribution.
@@ 3.2 Spatial Distribution Modeling
where $\mathcal{F}(\cdot,\cdot)$ is the joint operation of RoIAlign He et al. (2017) and two shared fully-connected layers, and $\mathbf{X}$ is the feature maps produced by the backbone. As a result, each proposal $P^{i}_{j}$ corresponds to a $D$-dimensional feature vector $F^{i}_{j}$. Then, we adopt the regressor $\mathcal{R}(\cdot)$ of the DA head to predict proposal offsets for further localization and update the features, which is formulated as
$\mathbf{P}^{i*}=\operatorname{Trans}(\mathbf{P}^{i},\mathcal{R}(\mathbf{F}^{i}))\in\mathbb{R}^{N^{i}\times 4},$ (2)
$\mathbf{F}^{i*}=\mathcal{F}(\mathbf{P}^{i*},\mathbf{X})\in\mathbb{R}^{N^{i}\times D},$ (3)
where $\operatorname{Trans(\cdot,\cdot)}$ is a function that translates predicted offsets to proposals Ren et al. (2015). Following Liu et al. (2022), we utilize classification scores to measure the possibilities of object locations. Therefore, the classifier $\mathcal{C}(\cdot)$ is used to score the proposals $\mathbf{F}^{i*}$, which is defined as
$\mathbf{S}^{i}=\mathcal{C}(\mathbf{F}^{i*})\in\mathbb{R}^{N^{i}\times(L+1)},$ (4)
$S^{i}=\operatorname{LookUp}(\mathbf{S}^{i},l^{i})=[s^{i}_{1},s^{i}_{2},...,s^{i}_{N^{i}}]\in\mathbb{R}^{N^{i}},$ (5)
@@
where $\operatorname{Softmax}(\cdot,\cdot)$ is the Softmax function to obtain normalized weights (sum up to $1$) and $T$ is the temperature coefficient Hinton et al. (2015) to control its sharpness. Finally, we model the spatial distribution of $\mathbf{P}^{i*}$ as a four-dimensional Gaussian distribution by directly calculating its parameters (i.e., mean $\mu^{i}$ and standard deviation $\sigma^{i}$) in a weighting manner, which can be formulated as
$\mu^{i}=\sum_{j=1}^{N^{i}}w^{i}_{j}*P^{i*}_{j}\in\mathbb{R}^{4},$ (7)
$\sigma^{i}=\sqrt{\sum_{j=1}^{N^{i}}w^{i}_{j}*(P^{i}_{j}-\mu^{i})^{2}}\in\mathbb{R}^{4},$ (8)
where we assume that each dimension of this Gaussian distribution is uncorrelated so that its standard deviation can be formulated as a four-dimensional vector. Note that $w^{i}_{j}$ is already normalized so dividing the sum is unnecessary. This assumption simplifies the modeling problem, makes the method more computationally efficient, and is not detrimental to performance. Therefore, it has been widely adopted in previous works, such as KL Loss He et al. (2019), GFL Li et al. (2020b), and Gaussian YOLOv3 Choi et al. (2019).

### v3

@@ Abstract
Large-scale well-annotated datasets are of great importance for training an effective object detector. However, obtaining accurate bounding box annotations is laborious and demanding. Unfortunately, the resultant noisy bounding boxes could cause corrupt supervision signals and thus diminish detection performance. Motivated by the observation that the real ground-truth is usually situated in the aggregation region of the proposals assigned to a noisy ground-truth, we propose DIStribution-aware CalibratiOn (DISCO) to model the spatial distribution of proposals for calibrating supervision signals. In DISCO, spatial distribution modeling is performed to statistically extract the potential locations of objects. Based on the modeled distribution, three distribution-aware techniques, i.e., distribution-aware proposal augmentation (DA-Aug), distribution-aware box refinement (DA-Ref), and distribution-aware confidence estimation (DA-Est), are developed to improve classification, localization, and interpretability, respectively. Extensive experiments demonstrate that DISCO can achieve SOTA performance in this task, especially at high noise levels. Code is available at https://github.com/Correr-Zhou/DISCO.
@@ 1 Introduction
Object detection has made substantial progress in recent years [Ren et al.(2015)Ren, He, Girshick, and Sun, Lin et al.(2017)Lin, Goyal, Girshick, He, and Dollár, Carion et al.(2020)Carion, Massa, Synnaeve, Usunier, Kirillov, and Zagoruyko, Sun et al.(2021)Sun, Zhang, Jiang, Kong, Xu, Zhan, Tomizuka, Li, Yuan, Wang, et al., Zou et al.(2023)Zou, Chen, Shi, Guo, and Ye], which is largely attributed to the utilization of large-scale well-annotated datasets [Everingham et al.(2010)Everingham, Van Gool, Williams, Winn, and Zisserman, Lin et al.(2014)Lin, Maire, Belongie, Hays, Perona, Ramanan, Dollár, and Zitnick]. However, obtaining accurate bounding box annotations is labor-intensive and demanding, especially for some real-world scenarios [Luo et al.(2021)Luo, Chen, Zhou, Lin, and Heng, Chai et al.(2023)Chai, Luo, Lin, Heng, and Chen, Michaelis et al.(2019)Michaelis, Mitzkus, Geirhos, Rusak, Bringmann, Ecker, Bethge, and Brendel, Mao et al.(2022)Mao, Shi, Wang, and Li]. Specifically, inherent ambiguities of bounding boxes are often caused by object occlusion or unclear boundaries [He et al.(2019)He, Zhu, Wang, Savvides, and Zhang]. Moreover, insufficient domain expertise and the strenuous workload can also lead to low-quality labeling of bounding boxes [Liu et al.(2022)Liu, Wang, Lu, Cao, and Zhang]. In practical deployments and applications, vanilla object detectors will inevitably suffer from such noisy bounding box annotations. Therefore, it is of scientific interest to explore how to tackle noisy bounding boxes in object detection.
@@
Figure 1: (a) Trait comparison of existing solutions and our DISCO. Their learning behaviors for one single border of bounding boxes are presented above. (b) Proposal aggregation in object detection with noisy bounding boxes. The real ground-truth is usually situated in the aggregation region of the proposals assigned to a noisy ground-truth.
Due to the degenerated supervision introduced by noisy annotations, object detection with noisy bounding boxes remains a challenging problem. Obviously, such corrupt supervision signals could weaken the localization precision of object detectors. Besides, although the classification accuracy is less affected [Liu et al.(2022)Liu, Wang, Lu, Cao, and Zhang], noisy bounding boxes do introduce biased category features during training, which reduces the generalization capability of classification. Notably, there is also a significant concern about the lack of interpretability for box predictions, especially considering the influence of noisy bounding box annotations.
Encountering the above challenges, existing solutions still exhibit drawbacks (see Figure 1(a)). Several methods are dedicated to correcting noisy bounding boxes [Liu et al.(2022)Liu, Wang, Lu, Cao, and Zhang, Li et al.(2020a)Li, Xiong, Socher, and Hoi, Xu et al.(2021)Xu, Zhu, Yang, and Wu], aiming to mitigate the effect of noise. However, their performance gains are constrained by heuristic box correction approaches, and the detector cannot identify which bounding boxes are inaccurately predicted. An alternative series of methods focuses on equipping the model with the ability to estimate the confidence of predicted bounding boxes [He et al.(2019)He, Zhu, Wang, Savvides, and Zhang, Li et al.(2020b)Li, Wang, Wu, Chen, Hu, Li, Tang, and Yang, Choi et al.(2019)Choi, Chun, Kim, and Lee], by which the detector can be more robust against noisy box annotations. Despite these efforts, it is unfortunate that the detector is still plagued by flawed supervision during training.
@@
Essentially, the corrupt supervision signals should be blamed for the above issues. In this work, we expect to answer the following question: How to properly calibrate the corrupt supervision signals? As shown in Figure 1(b), we observed that the real ground-truth is usually situated in the aggregation region of the proposals assigned to a noisy ground-truth, showing that the spatial distribution of proposals can act as a statistical prior for the potential locations of objects. Thus, we propose DIStribution-aware CalibratiOn (DISCO) to model the spatial distribution of proposals for calibrating supervision signals. For each group of the assigned proposals, we perform spatial distribution modeling, statistically extracting potential locations of objects. Based on the modeled distribution, we develop three distribution-aware techniques to improve classification, localization, and interpretability, respectively: 1) Distribution-aware proposal augmentation (DA-Aug): Additional proposals are generated from the distribution to enrich category features, and then the proposal with the highest classification score is collected to boost classification performance; 2) Distribution-aware box refinement (DA-Ref): With a non-linear weighting strategy, noisy ground-truth is fused with the distribution into a refined ground-truth to achieve superior bounding box regression; 3) Distribution-aware confidence estimation (DA-Est): An extra estimator is integrated into the detection head, with the distribution variance elegantly acting as its supervision, to estimate the confidence of box predictions. Our experiments show that DISCO can attain SOTA performance on large-scale noisy detection datasets, especially at high noise levels.
Our main contributions are listed as follows: (1) Motivated by the observation about proposal aggregation, we propose DISCO to calibrate supervision signals with spatial distribution modeling. (2) To improve classification, localization, and interpretability, we introduce three techniques (i.e., DA-Aug, DA-Ref, and DA-Est) to collaborate with the modeled distribution in a distribution-aware manner. (3) Comprehensive experiments show that DISCO can attain SOTA performance and achieve satisfactory interpretability for box predictions.
@@ 2 Related Works
Object Detection. The goal of object detection is to recognize what objects are present and where they are situated. Faster-RCNN [Ren et al.(2015)Ren, He, Girshick, and Sun] is a classic detection framework with a two-stage strategy, and is widely adopted and improved in subsequent works [Cai and Vasconcelos(2018), Pang et al.(2019)Pang, Chen, Shi, Feng, Ouyang, and Lin, Sun et al.(2021)Sun, Zhang, Jiang, Kong, Xu, Zhan, Tomizuka, Li, Yuan, Wang, et al.]. Moreover, RetinaNet [Lin et al.(2017)Lin, Goyal, Girshick, He, and Dollár], YOLO [Redmon et al.(2016)Redmon, Divvala, Girshick, and Farhadi], and CenterNet [Zhou et al.(2019)Zhou, Wang, and Krähenbühl] delve into strengthening the performance of one-stage detectors. Recently, transformer-based detectors [Carion et al.(2020)Carion, Massa, Synnaeve, Usunier, Kirillov, and Zagoruyko, Zhu et al.(2020)Zhu, Su, Lu, Li, Wang, and Dai, Zhang et al.(2022)Zhang, Li, Liu, Zhang, Su, Zhu, Ni, and Shum] also attracted the attention of the community, which conducts object detection in an end-to-end fashion. Trained with accurate box annotations, object detectors can achieve satisfactory performance. However, object detection with noisy bounding boxes remains an under-explored subproblem.
@@
Object Detection with Noisy Annotations. Noisy annotations of an object detection dataset comprise noisy category labels and noisy bounding boxes. Previous works [Chadwick and Newman(2019), Li et al.(2020a)Li, Xiong, Socher, and Hoi, Xu et al.(2021)Xu, Zhu, Yang, and Wu] have jointly handled these two types of noisy annotations. Unlike this setting, we focus on training an object detector with noisy bounding boxes, since box noise is more common and challenging in realistic scenarios [Liu et al.(2022)Liu, Wang, Lu, Cao, and Zhang]. The SOTA method for this tough task is called OA-MIL [Liu et al.(2022)Liu, Wang, Lu, Cao, and Zhang], which adopts a multi-instance learning (MIL) framework at the object level to correct bounding boxes. Besides, some approaches aiming to boost the robustness of detectors, such as KL Loss [He et al.(2019)He, Zhu, Wang, Savvides, and Zhang], can also contribute to performance improvement in this task.
@@
Weakly Supervised Object Detection. Weakly supervised object detection (WSOD) is also a relevant task, where only image-level labels can be accessed to train an object detector. The mainstream solution is to treat WSOD as a MIL problem [Bilen and Vedaldi(2016), Wan et al.(2018)Wan, Wei, Jiao, Han, and Ye, Chen et al.(2020)Chen, Fu, Jiang, Chen, and Hua, Tang et al.(2017)Tang, Wang, Bai, and Liu], where each training image is constructed as a bag of instances. To handle the non-convex optimization of MIL, spatial regularization [Diba et al.(2017)Diba, Sharma, Pazandeh, Pirsiavash, and Van Gool, Wan et al.(2018)Wan, Wei, Jiao, Han, and Ye], optimization strategy [Tang et al.(2017)Tang, Wang, Bai, and Liu, Wan et al.(2019)Wan, Liu, Ke, Ji, Jiao, and Ye], and context information [Kantorov et al.(2016)Kantorov, Oquab, Cho, and Laptev, Wei et al.(2018)Wei, Shen, Cheng, Shi, Xiong, Feng, and Huang] are introduced to attain better convergence. Moreover, it is worth noting that SD-LocNet [Zhang et al.(2019a)Zhang, Yang, and Feng] contributes a self-directed optimization strategy to handle object instances with noisy initial locations. Unfortunately, WSOD always results in relatively inaccurate box predictions due to the lack of fine-grained supervision. Effective methods for object detection with noisy bounding boxes can contribute to further refining these box predictions.
Figure 2: Training pipeline with DISCO. In DISCO, spatial distribution modeling (Section 3.1) is performed firstly, followed by three distribution-aware techniques, i.e., DA-Aug (Section 3.2), DA-Ref (Section 3.3), and DA-Est (Section 3.4), to collaborate with the modeled distribution. Note that we additionally integrate an estimator into the vanilla detection head to construct the distribution-aware head (DA head) for the implementation of DISCO.
@@ 3 Methodology
DISCO is a training-time approach designed for two-stage detectors. In a training iteration, it is performed twice using the distribution-aware head (DA head) with the assigned proposals as input (see Figure 2). Two times of DISCO follow the same process and have only subtle differences for different purposes. The first time aims to yield a refined ground-truth for proposal re-assignment, by which better-matching proposals can be obtained. The second time aims to produce spatial distributions of proposals acting as superior supervision. In the following, we start by describing spatial distribution modeling in Section 3.1. Then, we will introduce DA-Aug, DA-Ref, and DA-Est in Section 3.2, 3.3 and 3.4, respectively, in which we will also detail the differences between these two times of DISCO.
@@ 3.1 Spatial Distribution Modeling
where $\mathcal{F}(\cdot,\cdot)$ is the joint operation of RoIAlign [He et al.(2017)He, Gkioxari, Dollár, and Girshick] and two shared fully-connected layers, and $\mathbf{X}$ is the feature maps produced by the backbone. Each proposal $P^{i}_{j}$ corresponds to a $D$-dimensional feature vector $F^{i}_{j}$. Then, we adopt the regressor $\mathcal{R}(\cdot)$ of the DA head to predict proposal offsets for further localization and update the features, which is formulated as
$\mathbf{P}^{i*}=\operatorname{Trans}(\mathbf{P}^{i},\mathcal{R}(\mathbf{F}^{i}))\in\mathbb{R}^{N^{i}\times 4},\quad\mathbf{F}^{i*}=\mathcal{F}(\mathbf{P}^{i*},\mathbf{X})\in\mathbb{R}^{N^{i}\times D},$ (2)
where $\operatorname{Trans(\cdot,\cdot)}$ is a function that translates predicted offsets to proposals [Ren et al.(2015)Ren, He, Girshick, and Sun]. Following [Liu et al.(2022)Liu, Wang, Lu, Cao, and Zhang], we utilize classification scores to measure the possibilities of object locations. Therefore, the classifier $\mathcal{C}(\cdot)$ is used to score the proposals $\mathbf{F}^{i*}$, which is defined as
$\mathbf{S}^{i}=\mathcal{C}(\mathbf{F}^{i*})\in\mathbb{R}^{N^{i}\times(L+1)},\quad S^{i}=\operatorname{LookUp}(\mathbf{S}^{i},l^{i})=[s^{i}_{1},s^{i}_{2},...,s^{i}_{N^{i}}]\in\mathbb{R}^{N^{i}},$ (3)
@@
where $\operatorname{Softmax}(\cdot,\cdot)$ is the Softmax function to obtain normalized weights (sum up to $1$) and $T$ is the temperature coefficient [Hinton et al.(2015)Hinton, Vinyals, and Dean]. Finally, we model the spatial distribution of $\mathbf{P}^{i*}$ as a four-dimensional Gaussian distribution by directly calculating its parameters (i.e., mean $\mu^{i}$ and standard deviation $\sigma^{i}$) in a weighting manner, which can be formulated as
$\mu^{i}=\sum_{j=1}^{N^{i}}w^{i}_{j}*P^{i*}_{j}\in\mathbb{R}^{4},\quad\sigma^{i}=\sqrt{\sum_{j=1}^{N^{i}}w^{i}_{j}*(P^{i}_{j}-\mu^{i})^{2}}\in\mathbb{R}^{4},$ (5)
where we assume that each dimension of this Gaussian distribution is uncorrelated so that its standard deviation can be formulated as a four-dimensional vector. This assumption has also been widely adopted in previous works [He et al.(2019)He, Zhu, Wang, Savvides, and Zhang, Li et al.(2020b)Li, Wang, Wu, Chen, Hu, Li, Tang, and Yang, Choi et al.(2019)Choi, Chun, Kim, and Lee]. as it can simplify the modeling problem, make the method more computationally efficient, and is not detrimental to performance.
Figure 3: Illustration of spatial distribution modeling. For clarity, we present the process in the view of the whole box and one single border. Note that the length of the vertical line indicates its weight. Here the refined ground-truth is essentially a spatial distribution.